# TalkingEyes: Pluralistic Speech-Driven 3D Eye Gaze Animation

### v1

@@ V-C3 Qualitative Comparison
We visually compare our method with the four baselines in Fig. 6. It could be observed that, “Mean” can only generate stationary eye gaze motion, “Random” and “Similar Sample” tend to produce disorganized and speech-unrelated eyeball movement, and “VAE” results in motion that is nearly stationary. We strongly recommend readers to watch the supplementary video for more intuitive comparison. We also conduct a visual comparison with three recent speech-driven 3D facial animation methods, namely FaceFormer [15], CodeTalker [16] and Learn2Talk [19]. In Fig. 7, we illustrate three typical frames of synthesized facial animations that speak at specific syllables. It’s obvious that our method can generate the most vivid animation with the animated eye gaze in tune with the speech. In the supplementary video, we also show the visual comparison with an early speech-driven eye gaze animation method [32].
@@
[…]
@@ V-D2 Items in Training Losses
V-D3 Number of Codebook Items
[…]
@@ V-D3 Number of Codebook Items
[…]
@@
Figure 9: Visual comparisons of 3D eye gaze motions and 3D facial motions reconstructed from input video by using Wang et al.’s method [40] and our method.
@@ V-E Evaluation of LightGazeFit
We compare our LightGazeFit with a 3D eye gaze fitting method, Wang et al.’s method [40], and a deep-learning-based gaze estimation method, Gaze360 [49].
@@
As shown in Fig. 9, our method exhibits competitive performance in 3D eye gaze reconstruction, matching that of Wang et al.’s method [40], across both front view and side views. Our method employs only the pupil center as the fitting constraint, omitting the need for precise segmentation of the iris and pupil, as required by Wang et al.’s method. In this qualitative comparison, the input videos and the 3D reconstruction results of Wang et al.’s method are fetched from the supplementary video in [40], since the code of Wang et al.’s method has not been publicly released yet.
@@ V-E2 Comparison with Gaze360
Gaze360 [49] trains a deep neural network to predict 3D eye gaze direction from an input facial image. We conduct a quantitative comparison with Gaze360 and report the results in Tab. VIII. Specifically, we construct a evaluation dataset by manually annotating the 2D pupil centers in the facial images as ground truth. Then the 3D pupil centers generated by Gaze360 and our LightGazeFit are projected to the image plane and then compared with ground truth to measure the $L_{2}$ errors. Tab. VIII clearly show that our LightGazeFit achieves smaller prediction errors. Fig. 10 shows the visual comparison. There exists a deviation from the actual angle of eyeball rotation in Gaze360, indicating that the gaze direction produced by Gaze360 can’t be directly converted into the eye pose parameters in FLAME.
Method $L_{2}$
Gaze360 2.73
LightGazeFit (ours) 1.64
@@
Figure 10: Visual comparisons of 3D eye gaze motions estimated from input facial images by using Gaze360 [49] and our method.
@@ VI Conclusion and Limitations
In this paper, we study the speech-driven animation of an important facial component, eye gaze, which has been overlooked by the recent research. Through constructing the audio-gaze 3D dataset TKED and the explicit modeling of the translation from speech to 3D eye gaze motion, our proposed TalkingEyes can synthesize eye gaze motion, eye blinks, head motion and facial motion collectively from speech. Thanks to the cascade of VAE and VQVAE in learning two separate latent spaces, TalkingEyes can generate pluralistic and natural eye gaze motions and head motions.
One limitation of our method is that, currently the generation of eye gaze motion is not controlled by emotional labels. In addition to speech, eye gaze is weakly associated with human emotion. Implementing emotional control or constraint can help address the ill-posed problem in speech-driven 3D eye gaze animation. Another limitation is that the eye gaze and mouth movement are now separately generated. The coherence between these two parts can be enhanced to yield more coordinated and holistic motions. In future, we plan to continue improving our method along the above two directions.

### v2

@@ V-C3 Qualitative Comparison
Visual comparisons between our method and the four baselines are provided in the supplementary video. The comparison results demonstrate that, “Mean” can only generate stationary eye gaze motion, “Random” and “Similar Sample” tend to produce disorganized and speech-unrelated eyeball movement, and “VAE” yields nearly stationary motion. In contrast, our method produces natural eye gaze motion in tune with the speech. We also conduct visual comparisons with three recent speech-driven 3D facial animation methods in the supplementary video, namely FaceFormer [FaceFormer2022], CodeTalker [CodeTalker2023] and Learn2Talk [Learn2Talk2025]. In Fig. 6, we illustrate three typical frames of synthesized facial animations that speak at specific syllables. It’s obvious that our method generates the most vivid animation with the animated eye gaze that is synchronized with the speech. Finally, the supplementary video provides a visual comparison with an early speech-driven eye gaze animation method [LiveSpeehEyeTVCG2012].
@@
Method $\text{Div.}^{h}$ $\text{Div.}^{e}$ $\text{Corr.}^{h}$ $\text{Corr.}^{e}$ $L_{2}^{h}$ $L_{2}^{e}$
@@ V-D2 Items in Training Losses
Method $\text{Div.}^{h}$ $\text{Div.}^{e}$ $\text{Corr.}^{h}$ $\text{Corr.}^{e}$ $L_{2}^{h}$ $L_{2}^{e}$
@@
Num. $\text{Div.}^{h}$ $\text{Div.}^{e}$ $\text{Corr.}^{h}$ $\text{Corr.}^{e}$ $L_{2}^{h}$ $L_{2}^{e}$
@@
V-D3 Number of Codebook Items
Figure 8: Visual comparisons of 3D eye gaze motions and 3D facial motions reconstructed from input video by using Wang et al.’s method [wang2016realtime] and our method.
@@ V-E Evaluation of LightGazeFit
We compare our LightGazeFit with a 3D eye gaze fitting method, Wang et al.’s method [wang2016realtime], and a deep-learning-based gaze estimation method, Gaze360 [kellnhofer2019gaze360].
@@
As shown in Fig. 8, our method exhibits competitive performance in 3D eye gaze reconstruction, matching that of Wang et al.’s method [wang2016realtime], across both front view and side views. Our method employs only the pupil center as the fitting constraint, omitting the need for precise segmentation of the iris and pupil, as required by Wang et al.’s method. In this qualitative comparison, the input videos and the 3D reconstruction results of Wang et al.’s method are fetched from the supplementary video in [wang2016realtime], since the code of Wang et al.’s method has not been publicly released yet.
@@ V-E2 Comparison with Gaze360
We quantitatively compare our method with Gaze360 [kellnhofer2019gaze360], which trains a deep neural network to predict 3D eye gaze direction from an input facial image. The results are presented in Tab. VIII. To construct the evaluation dataset, we randomly selected 30 high-resolution videos (1080P) from HDTF [zhang2021flow] where the pupils are clearly visible. The ground truth 2D pupil centers in these video frames were obtained by using MediaPipe [MediaPipe2019, mediapipe] as the pupil center detector. We then degraded these high-resolution videos to match the resolution (224x224) and quality of VoxCeleb2 [chung2018voxceleb2], by applying downsampling and blur operations. This process created a realistic test bed for our LightGazeFit, which is designed to process low-resolution video. Both Gaze360 and LightGazeFit were used to process the test videos, and the resulting 3D pupil centers are projected to the image plane to measure $L_{2}$ error and gaze velocity. The $L_{2}$ error is defined as the Euclidean distance between the predicted 2D pupil centers and the ground truth, while the gaze velocity is defined as the speed of eyeball movement between adjacent frames, calculated as the 2D movement distance divided by the inter-frame interval of 40 ms.
Figure 9: Visual comparisons of 3D eye gaze motions estimated from input facial images by using Gaze360 [kellnhofer2019gaze360] and our method.
Method $L_{2}$ erorr↓ Velocity
Ground Truth 0.0000 16.2188
Gaze360 1.4041 20.9604
LightGazeFit (ours) 0.7311 18.5380
@@
As clearly shown in Tab. VIII, our LightGazeFit achieves superior performance, yielding both a lower prediction error and a gaze velocity that more closely matches the ground truth. Fig. 9 shows the visual comparison. There exists a deviation from the actual angle of eyeball rotation in Gaze360, indicating that the gaze direction produced by Gaze360 can’t be directly converted into the eye pose parameters in FLAME.
@@ VI Conclusion and Limitations
In this paper, we study the speech-driven animation of an important facial component, eye gaze, which has been overlooked by the recent research. Through constructing the audio-gaze 3D dataset TKED and the explicit modeling of the translation from speech to 3D eye gaze motion, our proposed TalkingEyes can synthesize eye gaze motion, eye blinks, head motion and facial motion collectively from speech. Thanks to VAE and VQVAE in learning two separate latent spaces, TalkingEyes can generate pluralistic and natural eye gaze motions and head motions.
Our method has several limitations that we plan to address in future work. 1) The exclusive reliance on speech input prevents our method from reacting to visual elements in a scene, such as interlocutors [S3TOG2024, migCanalesJJ23] or salient objects [SaliencyGazeTVCG2024]. This capability is crucial for fully interactive applications. A promising direction for our future work is to integrate visual inputs, such as the 3D locations of objects and interlocutors, to enable both goal-directed and speech-driven gaze fixation and enhance the model’s applicability in interactive virtual environments. 2) Additionally, our method can’t produce saccadic eye movements due to the 25 fps framerate of the training dataset. The resulting interval of 40 ms is too coarse to capture the fast dynamics of saccades. Addressing this limitation would require a dataset with a higher temporal resolution, which is a priority for future work. 3) Different individuals may exhibit different eye and head motion patterns for identical audio input due to factors like personal habit or cultural background. While our method employs a diverse generation strategy to address this variability, it does not support personalized generation for individual users. In future work, we will add speaker identity labels to TKED and use them as conditions during training to learn user-specific styles. 4) Finally, to further improve the physical realism of generated eye gaze in TKED, we will pursue a 3D eye gaze fitting method that incorporates biological constraints [misslisch1998neural] and $L_{0}$ gradient constraints [L02014, L02016, PointsL0].